You have found an asset with an arithmetic average return of 13.40 percent and a geometric average return of 10.40 percent. Your observation period is 30 years. What is your best estimate of the return of the asset over the next 5 years? 10 years? 20 years? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)
Answer #1

Return on assets = Arithmetic average return * (N - T) / (N - 1) + Geometric average * (T - 1) / (N - 1)

Where -

N is Observation period = 30 Years

T is the period for which estimate of return should be calculated = 5 years or 10 years or 20 years

5 Years = {13.40 * (30 - 5)} / (30 - 1) + 10.40 * (5 - 1) / (30 - 1)

= 0.11551724 + 0.01434482

= 0.1299 or 12.99%

The best estimate of return on asset over the next 5 years is 12.99%.

10 Years = {13.40 * (30 - 10)} / (30 - 1) + 10.40 * (10 - 1) / (30 - 1)

= 0.092413793 + 0.032275862

= 0.1247 or 12.47%

The best estimate of return on asset over the next 10 years is 12.47%.

20 Years = {13.40 * (30 - 20)} / (30 - 1) + 10.40 * (20 - 1) / (30 - 1)

= 0.04620689 + 0.06813793

= 0.1143 or 11.43%

The best estimate of return on asset over the next 20 years is 11.43%.